### v3

# A Survey on Video Prediction: From Deterministic to Generative Approaches

Abstract
Video prediction, a fundamental task in computer vision, aims to enable models to generate sequences of future frames based on existing video content. This task has garnered widespread application across various domains. In this paper, we comprehensively survey both historical and contemporary works in this field, encompassing the most widely used datasets and algorithms. Our survey scrutinizes the challenges and the evolving landscape of video prediction within the realm of computer vision. We propose a novel taxonomy centered on the stochastic nature of video prediction algorithms. This taxonomy accentuates the gradual transition from deterministic to generative prediction methodologies, underlining significant advancements and shifts in approach.
1 Introduction
The goal of the video prediction task is to forecast future frames given a sequence of historical frames. It is typically classified as a low-level computer vision task because it focuses on directly processing pixel values. Despite this classification, it diverges from other low-level tasks by implicitly demanding a more complex understanding of scene dynamics and temporal coherence, which is often characteristic of high-level vision tasks. The challenge lies in designing models that can achieve this balance efficiently, using a moderate number of parameters to minimize inference latency and resource consumption, thereby making video prediction suitable for real-world applications. This unique position of video prediction demonstrates its integral role in bridging the gap between low-level perception and high-level understanding within computer vision.
Current video prediction algorithms can generally be divided into two categories. One category entails flowing pixels from a reference frame (typically the last observed frame) to construct future frames. However, this group of methods inherently confronts difficulties in modeling objects’ appearance and disappearance (birth and death) within the scene. A shared feature of these methods is that they can make highly accurate predictions in the short term, but their accuracy diminishes over longer periods. The other category encompasses methods that involve generating new frames from scratch. Although these approaches offer the promise of capturing the birth-and-death phenomena in object dynamics, they predominantly focus on modeling pixel-level distributions. As a result, they often lack an integrated understanding of the underlying real-world context, which is crucial for creative prediction capabilities.
Figure 1: The overview of video prediction algorithms taxonomy
Recognizing these limitations, we propose a novel taxonomy, visually summarized in Figure 1, focusing on the stochasticity of algorithms. In Section 4, we introduce deterministic algorithms that aim to perform pixel-level fitting based on deterministic target frames. However, due to pixel-level metrics encouraging models to average over multiple equally probable outcomes, they often produce blurred outputs. In Section 5, we discuss algorithms that attempt to imbue models with the ability to make stochastic predictions in motion. This includes methods that introduce stochastic variables or distributions into deterministic models and approaches that directly utilize probabilistic models. Such algorithms allow models to sample from motion distributions, encouraging the generation of predictions that deviate significantly from the target frame but remain reasonable. Given the limited creativity of existing video prediction algorithms for high-resolution natural video data, which struggle with predicting video frame sequences containing many birth-and-death phenomena, we introduce the generative video prediction task in Section 6, which prioritizes generating reasonable video frame sequences in extended temporal durations over pixel-level accuracy. We believe that the future of long-term video synthesis hinges on a synergistic integration of video prediction and generation techniques. This unified approach would combine the contextual constraints offered by prediction methods with the enhanced comprehension of generation methods, addressing challenges within a cohesive framework.
2 Video prediction
2.1 Problem definition
Video prediction involves forecasting future frames or sequences in a video based on the analysis of past frames. The primary goal is to develop models that can accurately anticipate the visual content and probable motion of subsequent frames in a video sequence. This can be formulated as a conditional generative modeling problem, where given a sequence of observed frames $X_{t_{1}:t_{2}}$, the goal is to predict the future frames $Y_{t_{2}+1:t_{3}}$.
$Y_{t_{2}+1:t_{3}}=X_{t_{1}:t_{2}}\cdot P(Y_{t_{2}+1:t_{3}}|X_{t_{1}:t_{2}})$ (1)
Here, $t_{1}$ represents the initial time step, $t_{2}$ is the final time step for observed frames, and $t_{3}$ denotes the last time step for predicting future frames. The challenge lies in learning a mapping function that captures the complex spatiotemporal dependencies within the video sequence.
2.2 Application realms
The applications of video prediction span a wide range of domains, showcasing its significance in different fields: This task is crucial in various fields as it enables machines to plan and respond to dynamic environments.
Autonomous driving.
Video prediction is vital for self-driving cars and drones to anticipate the movement of objects, pedestrians, and other vehicles, ensuring safe navigation. GAIA-1 [42] employs a unified world model, combining multi-modal large language models and diffusion procedures to predict control signals and future frames.
Robot navigation.
Video prediction is employed in guiding robots through dynamic environments, allowing them to plan their paths, move objects, and avoid obstacles effectively [25].
Cinema industry.
Video prediction is utilized in special effects, animation, and pre-visualization, aiding filmmakers in creating realistic scenes and enhancing the overall cinematic experience [57].
Meteorological community.
Video prediction plays a crucial role in weather forecasting, helping meteorologists simulate and predict atmospheric conditions, contributing to improved accuracy in weather predictions [69].
World model.
World models [36] provide a more general framework for simulating and predicting the behavior of complex systems. A world model is often used in reinforcement learning and robotics to enable agents to make informed decisions and take actions that lead to desired outcomes. To construct a world model, video prediction is an important learning objective [37, 38, 84].
2.3 Challenges
In the video prediction field, there exist longstanding challenges, including the need for algorithms that balance low-level pixel processing with high-level scene dynamics understanding, the inadequacy of perceptual and stochastic evaluation metrics, the difficulty in achieving long-term predictions, and limited high-resolution dataset quality for stochastic motion and birth-and-death phenomena. This section provides an overview of challenges in the field of video prediction, emphasizing the ongoing need for research and development to enhance the accuracy, reliability, and adaptability of predictive models.
High-level metrics.
Widely used low-level, deterministic metrics such as Mean Squared Error (MSE), Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM), etc., can only assess the pixel-wise accuracy of predicted pixels, neglecting the visual performance and uncertain nature of videos. As a direct consequence, models are prone to average over multiple plausible outcomes, often resulting in blurry predictions that lack sharpness. Researchers are increasingly looking towards alternatives, such as perceptual metrics (DeePSiM [22], LPIPS [104]), and stochastic metrics (IS [65], FID [39]), that promise to promote more visually appealing and reasonable predictions.
Long-term prediction.
While short-term video prediction has seen significant advancements, accurately predicting events over extended time horizons remains challenging because of long-term dependencies and complex interactions between objects in dynamic scenes. Due to the inadequacy of model scale in forming a comprehensive understanding of the real world, most existing video prediction models primarily model changes in pixel distribution. When faced with natural videos spanning extended periods, these models struggle to accurately predict object movements while maintaining visual quality. A promising approach for enhancing long-term prediction capabilities is the incorporation of high-level structure information [79].
Dataset Year Category # Videos # Clip Frames Resolution Extra Annotations
KTH Action [66] 2004 Human 2,391 95 160 $\times$ 120 Class
Caltech Pedestrian [20] 2009 Human 137 1,824 640 $\times$ 480 Bounding Box
HMDB51 [49] 2011 Human 6,766 93 414 $\times$ 404$\dagger$ Class
UCF101 [73] 2012 Human 13,320 187 320 $\times$ 240 Class
J-HMDB [46] 2013 Human 928 34 320 $\times$ 240 OF, Ins, HJ, Class
KITTI [29] 2013 Traffic 151 323 1242 $\times$ 375 OF, BBox, Sem, Ins, Depth
Penn Action [106] 2013 Human 2,326 70 480 $\times$ 270$\dagger$ Human Joint, Class
SJTU 4K [72] 2013 General 15 300 3840 $\times$ 2160 -
Sports-1M [48] 2014 Human 1,133,158 variable variable Class
Moving MNIST [74] 2015 Simulation 10,000 20 64 $\times$ 64 -
Cityscapes [16] 2016 Traffic 46 869 2048 $\times$ 1024 Semantic, Instance, Depth
YouTube-8M [1] 2016 General 8,200,000 variable variable Class
Robotic Pushing [24] 2016 Robot 59,000 25 640 $\times$ 512 Class
DAVIS17 [61] 2017 General 150 73 3840 $\times$ 2026$\dagger$ Semantic
Something-Something [31] 2017 Object 220,847 45 427 $\times$ 240$\dagger$ Text
ShapeStacks [32] 2018 Simulation 36,000 16 224 $\times$ 224 Semantic
SM-MNIST [18] 2018 Simulation customize customize 64 $\times$ 64 -
RoboNet [17] 2019 Robot 161,000 93 64 $\times$ 48 -
Vimeo-90K [98] 2019 General 91,701 7 448 $\times$ 256 -
WebVid [6] 2021 General 10,732,607 449 596 $\times$ 336 Text
X4K1000FPS [71] 2021 General 4,408 65 4096 $\times$ 2160 -
denotes the mean value. $\dagger$ denotes the median value.
Table 1: Summary of the most used video prediction datasets, including the total number of videos, frame count for each video clip, image resolution, and additional annotations, etc. (OF: Optical Flow, BBox: Bounding Box, Sem: Semantic, Ins: Instance, HJ: Human Joints)
Generalization.
The interplay between data volume and the complexity of models jointly determines the upper bounds of an algorithm’s performance. Despite the vastness of video data available on the internet, the scarcity of high-quality video datasets suitable for video prediction remains a limiting factor, and existing datasets also pose various challenges, such as simplistic data distribution, low resolution, small motion scales, etc. These issues make it difficult for video prediction models to handle high-resolution content and large motion scales, thus restricting their practical utility to diverse and unseen scenarios. Achieving high-resolution video prediction is a complex task demanding significant computational resources [9]. The intricacies involved make real-time applications challenging due to the substantial computational burden.
3 Datasets
To assess the performance of video prediction models, we not only analyze their architectural advancements but also consider the impact of the datasets used for training and testing. The advancement in video prediction models is greatly dependent on the diversity, quality, and characteristics of these datasets. A general observation is the varying suitability of datasets based on their dimensionality and size, where lower-dimensional datasets, typically with smaller data sizes, may suffer from limited generalizability. In contrast, higher-dimensional datasets provide a broader range of data, contributing to stronger generalization capabilities in the models. Table 1 offers an overview of the most widely used datasets in video prediction, highlighting their amounts of data and additional supervisory modalities, thereby providing a comprehensive picture of the current dataset landscape in this field. In cases where specific dataset details are not reported either in the original paper or the project page, we calculate the mean or median statistics to maintain consistency in our analysis.
4 Deterministic prediction
4.1 Raw pixel space
Early research in video prediction was centered on raw pixel space modeling, a method that is straightforward but computationally demanding. Consequently, it was primarily applied to small-scale, low-resolution datasets.
Recurrent networks.
The exploration of recurrent neural networks in video prediction is pioneered by PredNet [55], which draws inspiration from predictive coding in neuroscience and employs a recurrent convolutional network for the effective processing of video features. Building on this, PredRNN [87] introduces significant enhancements by modifying Long Short-term Memory (LSTM) with a dual memory structure, aiming for enhanced spatiotemporal modeling. Despite its advancements, it encounters challenges with gradient vanishing in video prediction tasks. Addressing such limitations, ConvLSTM [68] emerges as a pivotal model, ingeniously integrating LSTM with Convolutional Neural Network (CNN) to proficiently capture motion and spatiotemporal dynamics, a development that has significantly influenced subsequent video prediction models. Further advancing the field, E3d-LSTM [86] innovatively incorporates 3D convolutions into RNNs and introduces a gate-controlled self-attention module, thereby markedly improving long-term prediction capabilities in video prediction.
Convolutional networks.
Convolutional neural networks have been instrumental in the evolution of video prediction technology. Beginning with GDL [58], the field has seen significant advancements. Following this, PredCNN [97] establishes a new benchmark by outperforming its predecessor PredRNN [87] across various datasets. SDC-Net [63] introduces a novel approach by utilizing a high-resolution video frame prediction technique that effectively leverages past frames and optical flow. Building on these innovations, the recent introduction of SimVP [28] marks another milestone. This approach revisits the advancements made by ViT [21] and introduces a simplified CNN network, demonstrating that such a configuration can achieve comparable performance in video prediction. This progression underscores the rapid advancements in convolutional network applications in video prediction, highlighting a continuous trend toward more efficient and effective models.
4.2 Feature space
Predicting in raw pixel space often overburdens models by requiring them to reconstruct images from scratch, a task particularly challenging for high-resolution video datasets. This realization has led to a shift in focus among researchers. Rather than grappling with the complexities of pixel-level prediction, several studies have pivoted towards high-level feature predictions in feature space, such as optical flow and segmentation maps. These approaches offer a more efficient way of handling the intricacies of videos [60].
Optical flow prediction.
The field of video prediction has seen significant advancements through innovative approaches. Focusing on the enhancement of prediction quality, FVS [93] utilizes a comprehensive approach by incorporating supplementary information such as semantic maps, instance maps, and optical flow from input frame sequences. This method, while effective, introduces challenges due to increased data modalities and computational demands. In a related development, the research by OPT [94] draws inspiration from the successes in photo-realistic video interpolation. It presents an optimization framework specifically tailored for video prediction, which is grounded in the principles of an advanced interpolation algorithm [44]. Building on these concepts, DMVFN [43] represents a further evolution in this field. It extends the idea of dense voxel flow [54] and integrates a differentiable routing module. This addition is pivotal in enabling the model to more effectively capture and represent varying scales of motion, showcasing the continual progression and refinement in the methods of video prediction based on optical flow.
Future semantic segmentation.
Future semantic segmentation represents a progressive approach of video prediction, primarily focusing on predicting semantic maps for forthcoming video frames. This methodology diverges from traditional raw pixel forecasting, turning to semantic maps to narrow the prediction scope and enrich scene comprehension. In this context, the S2S model [56] stands as a groundbreaking end-to-end system. It processes RGB frames alongside their semantic maps, both as input and output. This integration not only advances future semantic segmentation but also elevates the task of video frame prediction, showcasing the distinct advantage of semantic-level forecasting. Building on this foundation, SADM [7] further innovates by amalgamating optical flow with semantic maps. This fusion leverages optical flow for motion tracking and semantic maps for appearance detailing, employing the former to warp input images and the latter to inpaint occluded areas.
5 Stochastic prediction
In the early stages, video prediction is primarily perceived as a low-level computer vision task, with a focus on employing deterministic algorithms to enhance pixel-level metrics such as MSE, PSNR, and SSIM. This approach, however, inherently limits the potential for creative output from these models by confining possible motion outcomes to a single, fixed result [60]. This often results in high pixel-level scores but at the cost of producing blurry images, which significantly diminishes the practical applicability of these algorithms. Recognizing this issue, there has been a paradigm shift in the field of video prediction, moving away from the reliance on short-term deterministic predictions towards embracing long-term stochastic predictions. This transition acknowledges that while stochastic prediction may yield results that diverge significantly from the ground truth, it plays a crucial role in fostering a more comprehensive understanding and enhancing the creative predictive capabilities concerning the evolution of video content.
5.1 Stochasticity modeling
Modeling the uncertain motion of the objects is commonly achieved by introducing stochastic distributions into deterministic models, or directly leveraging probabilistic models.
5.1.1 Stochastic distributions
In the early stage, VPN [47] uses CNNs for multiple predictions in videos based on pixel distributions and SV2P [4] enhances an action-conditioned model [24] with stochastic distribution estimation for videos. Shifting the focus to a more holistic view of video elements, PFP model [41] presents a probabilistic method for predicting semantic segmentation, depth map, and optical flow in videos simultaneously. Additionally, SRVP [26] utilizes Ordinary Differential Equations (ODEs), while PhyDNet [34] employs Partial Differential Equations (PDEs) to calculate stochastic distributions.
5.1.2 Probabilistic models
With the pioneering work of GDL [58] establishing the foundation, adversarial training has significantly advanced video prediction tasks in predicting uncertain object motions. Similarly, vRNN [12] and GHVAE [91] enhance VAEs through likelihood networks and hierarchical structures respectively, thereby offering another dimension to the ongoing evolution of stochastic prediction methodologies.
Recognizing that object motion is largely deterministic barring unforeseen events like collisions, SVG [18] models trajectory uncertainty using fixed and learnable priors, which effectively blends deterministic and probabilistic approaches. In a similar vein, but with a focus on enhancing temporal aspects, Retrospective Cycle GAN [50] introduces a sequence discriminator designed to detect fake frames. This idea of scrutinizing frame authenticity is further extended in DIGAN [103], where the emphasis shifts to a motion discriminator that concentrates on identifying unnatural motions.
To overcome pixel-level prediction challenges in stochastic models, several works introduce intermediate representations. S2S [56] and Vid2Vid [82] integrate adversarial training with future semantic segmentation. Additionally, TPK model [81] leverages a VAE to extract human pose information, and then a GAN for predicting future poses and frames.
It is worth noting that directly modeling stochastic distributions is prone to covering a broader predictive distribution but suffers from poor visual effects. In contrast, probabilistic models are capable of producing sharper results, yet they grapple with issues such as mode collapsing, training difficulties, and substantial computational overhead. Bridging these two methodologies, SAVP [51] combines stochastic distributions with adversarial training aiming to achieve broader predictive distributions with promising visual quality.
5.2 Disentangling components
Stochastic prediction algorithms primarily focus on the randomness in motion. This approach, however, often neglects the birth-and-death phenomena of objects in videos. Consequently, numerous studies isolate motion from other video elements or artificially manipulate its evolution, providing a clearer understanding of motion dynamics while simplifying the complexity of real-world scenarios.
5.2.1 Content and motion
Video prediction algorithms confront the challenging complexity of natural video sequences by emphasizing intricate image details. To this end, they focus on effectively modeling appearances through detailed local information, while simultaneously developing a comprehensive understanding of the dynamic global content inherent in videos. However, in applications such as robot navigation and autonomous driving, the importance of understanding object motion patterns supersedes the pursuit of visual aesthetics. This priority shift has encouraged the development of algorithms that emphasize predicting the movement of objects and distinguishing motion from appearance in videos. Early work CDNA [24] has set a precedent by explicitly predicting the motion of objects. It maintains invariant appearance characteristics, which is instrumental in applying the model to objects beyond those encountered during training.
MoCoGAN [77] automatically learns to disentangle motion from content in an unsupervised manner, while the approach of leveraging separate content and motion encoder pathways has also been widely used in various video prediction models. This idea of separating content and motion is further explored in LMC [52], which makes the motion encoder focused on motion prediction based on residual frames, while the content encoder extracts content features from the input frame sequence. MMVP [107] takes a different approach by employing only one image encoder to extract information, and then a two-stream network before the image decoder to separately handle motion prediction and appearance maintenance.
Addressing the stochastic nature of motion, AMC-GAN [45] models multiple plausible outcomes via adversarial training. Transitioning to a different approach, SLAMP [3] adopts a non-adversarial approach but focuses on learning stochastic variables for separate content and motion. Further advancing this field, LEO [88] and D-VDM [67] leverage diffusion models for more realistic disentanglement of content and motion, showcasing the latest advancements in this direction.
5.2.2 Foreground and background
In the realm of video prediction, the motion dynamics of foreground (objects) and background (scenes) often display stark contrasts. Foreground objects typically exhibit more intense movement, whereas scenes tend to remain relatively static. This distinction has steered research towards predicting the motions of these elements separately, offering a more nuanced understanding of video dynamics. A pivotal contribution in this area is DrNet [19], which specifically addresses scenarios where the background remains largely unchanged across video frames. Focusing on such videos, DrNet streamlines the prediction process by eliminating the need to learn complex scene dynamics. This model ingeniously decomposes images into the content and pose of objects. It then harnesses adversarial training techniques to develop a scene discriminator that evaluates whether two pose vectors belong to the same video sequence.
Human-centric.
Video prediction often centers around foreground motion, particularly when it involves intricate human movements. In such scenarios, a common assumption across various specialized datasets is the relatively static nature of the background, which is a typical characteristic in datasets focusing on detailed human motion. This has led to a significant focus in research on understanding and forecasting human poses to enhance the prediction of foreground motion. An example of this approach is seen in the work of DVGPC [11], which innovatively predicts skeleton motion sequences before transforming these sequences into pixel space using a skeleton-to-image transformer. Their method is an effective solution for linking the abstract representation of motion with the practical aspect of video prediction, demonstrating a nuanced understanding of the complexities involved in human-centric video prediction tasks.
Object-centric.
The field of Object-Centric Video Prediction (OCVP) focuses on video datasets particularly aimed at predicting the motion of objects. This concept is first introduced in the work of CVP [100], which lays the foundation for this specialized area of video prediction. SlotFormer [95] introduces transformer-based autoregressive models to learn representations of each object in video sequences. This innovation ensures consistent and accurate tracking of each object over time. The OKID model [15] represents a recent advancement. It uniquely decomposes videos into distinct elements, specifically the attributes and trajectory dynamics of moving objects, employing a Koopman operator. This approach highlights a more detailed method of analyzing object motion within video sequences, differentiating it from previous methodologies.
General.
Methods that concentrate on human poses or objects have demonstrated considerable promise in specific video datasets, yet they encounter limitations due to their reliance on pre-defined structures and struggle with adapting to variable backgrounds, which hampers their ability to generalize. This challenge is evident in their performance, which, while effective under certain conditions, falters when faced with dynamic background shifts, revealing a lack of the versatility required for wider applications. Bridging this gap, MOSO [75] emerges as a notable approach, identifying motion, scene, and object as the pivotal elements of a video. It delves deeper into content analysis by distinguishing between scene and object. Scene and object are considered as a further breakdown of content, where scene represents the background and object represents the foreground. MOSO’s innovative contribution lies in its two-stage network tailored for general video analysis. Initially, the MOSO-VQVAE model dissects video frames into token-level representations, honing its capabilities through a video reconstruction task. Subsequently, the model employs transformers in its second stage, tackling disparities in masked tokens. This strategic design equips the model to handle a variety of tasks at the token level, including video prediction, interpolation, unconditional video generation, etc.
5.3 Motion-controllable prediction
In the field of video prediction, a specialized research direction has emerged that focuses on the explicit control of motion. This approach is distinct in its emphasis on forecasting future object positions based on user-defined instructions, diverging from the conventional reliance on past motion trends. The central challenge in this domain lies in the synthesis of videos that adhere to these direct instructions while maintaining a natural and coherent flow, a task that demands a nuanced understanding of both user intent and the dynamics of motion within a video context. This challenge highlights the intricate balance between user control and automated prediction, marking a significant shift in how video prediction models are conceptualized and implemented.
Strokes.
There is no historical motion information to be used for video prediction from one still image, hence several methods have emerged allowing for interactive user control. iPOKE [8] introduces techniques where local interactive strokes and pokes enable users to deform objects in one still image to generate a sequence of video frames. These strokes indicate the user’s intended motion for the objects. Following this innovative pathway, Controllable-Cinemagraphs model [57] introduces a method to interactively control the animation of fluid elements. The advancements in the field underscore the growing importance of user-centric approaches in the realm of motion-controllable video prediction.
Instructions.
The integration of instructions of various modalities, including local strokes, sketches, and texts, is increasingly common in works aiming to capture user-specified motion trends. VideoComposer [83] synthesizes videos by combining text descriptions, hand-drawn strokes, and sketches. This approach respects textual, spatial, and temporal constraints, leveraging video latent diffusion models and motion vectors for explicit dynamic guidance. Essentially, It can generate videos that align with user-defined motion strokes and shape sketches. In a similar vein, DragNUWA [101] primarily leverages text for video content description and strokes for future motion control, enabling the prediction of customizable videos. These attempts advance the field of video prediction by broadening the range of user input modalities.
6 Generative prediction
In video analysis, the focus shifts towards algorithms designed for generative video prediction, especially when dealing with videos that exhibit stochastic birth-and-death events. These events introduce unpredictability as objects come into existence and vanish. These algorithms necessitate a profound understanding of the underlying physical principles governing the real world to tackle such complexities. Instead of relying on simplistic linear motion predictions extrapolated from historical frames, they embrace the challenge with sophisticated and imaginative modeling techniques. As a result, tasks like transforming a single static image into a dynamic video, often referred to as the image animation problem, emerge as promising candidates for applying generative video prediction techniques.
6.1 Unguided prediction
In computer vision, addressing the non-sequential nature of pixels poses a challenge in transformer architecture. Innovations like VQ-VAE [78] and VQGAN [23] combine autoregressive models and adversarial training strategies to tackle image quantization. Subsequently, VideoGPT [99] and LVM [5] use VQ-VAE and VQGAN, respectively, to construct models for natural video prediction.
Before the rise of Large Language Models (LLMs) [2], the transformative impact of transformers has been evident in time-series modeling. Video Transformer [89] applies transformers to create an autoregressive model, pioneering the application of transformer architectures in the context of video prediction. While this autoregressive model marked a successful attempt in the realm of natural video prediction, offering the potential for generatively predicting videos, it also inherited the consistent drawbacks of transformers, namely extensive training resource consumption and prolonged inference time. The Latent Video Transformer (LVT) [62] proposed a novel approach by considering the modeling of dynamics and predicting future features in an autoregressive manner within a latent space to release the computation burden.
The NÜWA [92] framework proposes a versatile 3D transformer encoder-decoder architecture, adaptable to various data modalities and tasks, further highlighting the potential of transformers in video prediction. NUWA-Infinity [53] expands on this with an innovative generation mechanism, with the ambitious goal of achieving infinite high-resolution video prediction. This progress signifies a continuous effort to unify generative tasks across different modalities, showcasing the evolving landscape of video prediction research.
Diffusion models [40] have become dominant in the image generation realm. The Latent Diffusion Model (LDM) [64] explores the generation capability in the latent space of images, significantly improving computational efficiency and reducing costs. LDM has paved the way for introducing diffusion models to the video prediction realm. The extensions of LDM for video prediction exhibit robust generative capabilities [80]. Furthermore, Video LDM [9] leverages pre-trained image models for video generation, offering multi-modal, high-resolution, long-term video prediction results. Similarly, SEINE [14] introduces a versatile video diffusion model, enabling the creation of transition sequences to generate longer videos from shorter clips. CVP model [70] introduces treating video as a continuous multi-dimensional process rather than a series of discrete frames can achieve impressive performance in complex human motion prediction.
Although diffusion models have almost become the most popular generative models, there are many researchers attempting to preserve the architecture of autoregressive models to simultaneously utilize the advantages of both mechanisms [90].
Despite the limitations in data availability and model scalability that hindered the early generative video prediction algorithms mentioned above, they laid the groundwork for integrating LLMs and diffusion models in the field of video prediction. Recently, Sora [10] and Kolors [76] leverage the strengths of both diffusion models and large language models to achieve more creative, general and scalable video prediction.
6.2 Text-guided prediction
Text-guided generative video prediction algorithms are designed to create a sequence of frames by integrating context frames and textual guidance. Recent examples include MMVG [27] and MAGVIT [102] addressing the novel Text-guided Video Completion (TVC) task. This task involves completing videos based on various conditions, including the first frame (video prediction task), the last frame (video rewind task), or a combination of both (video transition task), all of which are guided by textual instructions. These models utilize autoregressive encoder-decoder architectures, integrating text and frame features, resulting in a unified framework capable of handling multiple video synthesis tasks.
Recently, there has been a surge in research efforts aimed at harnessing additional information in conjunction with RGB images to fulfill the text-guided video completion task. LFDM [59] builds upon the latent diffusion models to synthesize optical flow sequences in latent space based on textual guidance. Seer [33] inflates Stable Diffusion [64] along the temporal axis, enabling the model to utilize natural language instructions and reference frames to predict multiple variations of future outcomes. Emu Video [30] takes a different approach by initially generating an image conditioned on textual guidance and subsequently extrapolating it into a video. Consequently, this method can be flexibly applied to generatively predict videos based on a given image with different textual inputs. DynamiCrafter [96] broadens the application of text-guided image animation to open-domain images. SparseCtrl [35] allows for sketch-to-video generation, depth-to-video generation, and video prediction with an expanded input range.
Some works attempt to provide a more precise understanding of the guidance of text and preserve the guidance information along the time dimension. MicroCinema [85] adopts a divide-and-conquer strategy to tackle appearance and temporal coherence challenges. This model employs a two-stage generation pipeline, which initially utilizes any existing text-to-image generator for the creation of the initial image, and subsequently introduces a dedicated text-guided video generation framework specifically designed for motion modeling. LivePhoto [13] introduces a framework that incorporates motion intensity as a supplementary factor to enhance control over desired motions. Additionally, it proposes a text re-weighting mechanism to emphasize motion descriptions, showcasing impressive performance on text-guided video prediction tasks. I2VGen-XL [105] utilizes static images for semantic and qualitative guidance, showcasing the diverse approaches in text-guided video prediction research.
7 Future research
Based on the overall development trends in the fields of Artificial Intelligence and Computer Vision, and considering the challenges in the Video Prediction domain discussed in Section 2.3, we contend that the trajectory of video prediction research should diverge from the traditional trends of data simplification and model lightweight in low-level tasks. The imperative should be to harness substantial computational resources and high-resolution, long-duration video data. Future research should prioritize the formulation of evaluation metrics that incentivize the generation of stochastic predictions, thus expanding the potential of model performance. The ultimate objective is to cultivate video prediction models with a profound comprehension of the inherent dynamics within videos. Such models could execute predictions over extended time horizons with high stochastic complexity in the real world.
8 Conclusion
In this survey, we have discussed various aspects of video prediction, covering key topics such as predominant datasets, evolving algorithms, current challenges, and future trends. We propose a novel taxonomy based on algorithm stochasticity, dividing video prediction into deterministic, stochastic, and generative prediction, showcasing a significant shift from low-level deterministic approaches toward high-level creative methodologies. This survey emphasizes the importance of balancing pixel-level accuracy with a deep understanding of complex scene dynamics in video prediction. Additionally, we delve into the intricacies of stochastic birth-and-death phenomena, advocating for enhanced evaluation metrics and the utilization of substantial computational resources and large-scale video datasets. These insights aim to guide future research directions in video prediction.
In conclusion, research in video prediction is of critical importance and has the potential to significantly impact various downstream applications. As the field advances, we anticipate the development of models that provide a more profound and nuanced grasp of nature in the real world. This evolution promises to enhance predictions’ accuracy, efficiency, and creativity, paving the way for novel applications and research. Such advancements will significantly contribute to the broader domains of computer vision and artificial intelligence.
Acknowledgments
This work is supported by Beijing Natural Science Foundation L231014.