### v4

# A Survey on Future Frame Synthesis: Bridging Deterministic and Generative Approaches

Abstract
Future Frame Synthesis (FFS) aims to enable models to generate sequences of future frames based on existing content. This task has garnered widespread application across various domains. In this paper, we comprehensively survey both historical and contemporary works in this field, encompassing widely used datasets and algorithms. Our survey scrutinizes the challenges and the evolving landscape of FFS within the realm of computer vision. We propose a novel taxonomy centered on the stochastic nature of related algorithms. This taxonomy emphasizes the gradual transition from deterministic to generative synthesis methodologies, highlighting significant advancements and shifts in approach.
1 Introduction
The goal of the future frame synthesis (FFS) task is to generate future frames based on a sequence of historical frames [90] or just a single context frame [122], with or without additional control signals. It is closely related to low-level computer vision processing techniques, particularly when synthesizing near frames [70, 50]. However, FFS diverges from other low-level tasks by implicitly demanding a more complex understanding of scene dynamics and temporal coherence, which is often characteristic of high-level vision tasks. The challenge lies in designing models that can achieve this balance efficiently, using a moderate number of parameters to minimize inference latency and resource consumption, thereby making FFS suitable for real-world applications. This unique position of FFS demonstrates its integral role in bridging the gap between low-level perception & prediction and high-level understanding & generation within computer vision.
Current future frame synthesis algorithms can generally be divided into two categories. One category entails referencing pixels from the current frames (typically the last observed frame) to construct future frames. However, this group of methods inherently confronts difficulties in modeling objects’ appearance and disappearance (birth and death) within the scene. A shared feature of these methods is that they can make highly accurate predictions in the short term, but their accuracy diminishes over longer periods. This kind of research is also commonly referred to as video prediction [76]. The other category encompasses methods that involve generating new frames from scratch. Although these approaches offer the promise of capturing the birth-and-death phenomena in object dynamics, they predominantly focus on modeling pixel-level distributions. As a result, they often lack an integrated understanding of the underlying real-world context, which is crucial for creative synthesis capabilities.
Before our work, two reviews in 2020 on video prediction [76, 79] cover a lot of early technical content. Recently, there have also been literature reviews on “text-to-video generative model” [69] and “long video generation” [65, 93]. Our survey mainly highlights the latest advancements and the connections between predictive and generative methodologies. We believe that the future of long-term FFS hinges on a synergistic integration of prediction and generation techniques. This unified approach would combine contextual constraints with enhanced comprehension, addressing challenges within a cohesive framework.
We propose a novel taxonomy, visually summarized in Figure 1, focusing on the stochasticity of methods. In Section 4, we introduce deterministic algorithms that aim to perform pixel-level fitting based on deterministic target frames. However, due to pixel-level metrics encouraging models to average over multiple equally probable outcomes, they often produce blurred outputs. In Section 5, we discuss algorithms that attempt to imbue models with the ability to make stochastic predictions in motion. This includes methods that introduce stochastic variables or distributions into deterministic models and approaches that directly utilize probabilistic models. Such algorithms allow models to sample from motion distributions, encouraging the generation of predictions that deviate significantly from the target frame but remain reasonable. Given the limited creativity of existing FFS algorithms for high-resolution natural video data, which struggle with creating future frame sequences containing many birth-and-death phenomena, we introduce the generative FFS task in Section 6, which prioritizes generating reasonable video frame sequences in extended temporal durations over pixel-level accuracy. In addition, some related content is placed in the Appendix, including the application scenarios of FFS and the detailed information the training datasets.
2 Future frame synthesis
2.1 Problem definition
The nature of FFS task involves forecasting future frames or sequences in a video based on the analysis of past frames. The primary goal is to develop models that can accurately anticipate the visual content and probable motion of subsequent frames in a video sequence. This can be formulated as a conditional generative modeling problem, where given a sequence of observed frames $X_{t_{1}:t_{2}}$, the goal is to synthesize the future frames $Y_{t_{2}+1:t_{3}}$.
Figure 1: The overview of FFS methods taxonomy
$Y_{t_{2}+1:t_{3}}=X_{t_{1}:t_{2}}\cdot P(Y_{t_{2}+1:t_{3}}|X_{t_{1}:t_{2}}).$ (1)
In Eq. 1, $t_{1}$ represents the initial time step, $t_{2}$ is the final time step for observed frames, and $t_{3}$ denotes the last time step for synthesizing future frames. The challenge lies in learning a mapping function that captures the complex spatiotemporal dependencies within the video sequence.
Many FFS algorithms also make use of additional information, which can be auxiliary data $A_{t_{1}:t_{2}}$ from videos, such as depth maps, landmarks, bounding boxes and segmentation maps, to help the model understand the video. It can also include human input control signals $C_{t_{2}+1:t_{3}}$, such as text instructions and strokes, with the goal of enabling the model to generate future video frame sequences based on specific future trajectory. Based on this consideration, we can extend the formula of the FFS task to a more comprehensive version as Eq. 2:
$\small Y_{t_{2}+1:t_{3}}=X_{t_{1}:t_{2}}\cdot P(Y_{t_{2}+1:t_{3}}|X_{t_{1}:t_{2}},A_{t_{1}:t_{2}},C_{t_{2}+1:t_{3}}).$ (2)
2.2 Challenges
There exist longstanding challenges in FFS field, including the need for algorithms that balance low-level pixel processing with high-level scene dynamics understanding, the inadequacy of perceptual and stochastic evaluation metrics, the difficulty in achieving long-term synthesis, and limited high-resolution dataset quality for stochastic motion and birth-and-death phenomena. This section provides an overview of challenges, emphasizing the ongoing need for research and development to enhance the accuracy, reliability, and adaptability of FFS algorithms.
High-level metrics.
The low-level metrics such as Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM), can only assess the pixel-wise accuracy of predicted pixels. To chasing these metrics, researchers typically train models utilizing either the $l_{1}$ or $l_{2}$ loss function of pixel space. The models are prone to average over multiple plausible outcomes, often resulting in blurry predictions know as the perception-distortion tradeoff [10]. Researchers are increasingly looking towards alternatives, such as perceptual metrics (DeePSiM [29], LPIPS [130]), and stochastic metrics (IS [82], FID [46]). These metrics may better align with human perception. At present, reseachers generally use Imagenet classifiers as feature extractors [55, 61] in comparing the generated results with the ground truth to optimize for perception metrics.
Long-term synthesis.
While short-term video prediction has seen significant advancements, accurately synthesizing events over extended time horizons remains challenging because of long-term dependencies and complex interactions between objects in dynamic scenes. Due to the inadequacy of model scale in forming a comprehensive understanding of the real world, most existing video synthesis models primarily model changes in pixel distribution. When faced with natural videos spanning extended periods, these models struggle to accurately predict object movements while maintaining visual quality. A promising approach for enhancing long-term synthesis capabilities is the incorporation of high-level structure information [99]. Understanding and utilizing higher-order information can enable the model to retain more key details, thereby achieving internal consistency within the video over longer time scales.
Generalization.
The interplay between data volume and the complexity of models jointly determines the upper bounds of an algorithm’s performance. Despite the vastness of video data available on the internet, the scarcity of high-quality video datasets suitable for video synthesis remains a limiting factor, and existing datasets also pose various challenges, such as simplistic data distribution, low resolution and small motion scales. These issues make it difficult for video synthesis models to handle high-resolution content and large motion scales, thus restricting their practical utility to diverse and unseen scenarios. Achieving high-resolution video synthesis is a complex task demanding significant computational resources [9]. The intricacies involved make real-time applications challenging due to the substantial computational burden.
3 Datasets
The advancement in video synthesis models is greatly dependent on the diversity, quality, and characteristics of training datasets. A general observation is the varying suitability of datasets based on their dimensionality and size, where lower-dimensional datasets, typically with smaller data sizes, may suffer from limited generalizability. In contrast, higher-dimensional datasets provide a broader range of data, contributing to stronger generalization capabilities in the models. Our Appendix offers an overview of the most widely used datasets [83, 26, 60, 89, 54, 36, 132, 88, 57, 90, 22, 1, 31, 77, 38, 39, 24, 23, 121, 6, 87, 126, 123, 12, 17, 15, 109] in video synthesis, highlighting their amounts of data and additional supervisory modalities, thereby providing a comprehensive picture of the current dataset landscape in this field.
4 Deterministic synthesis
4.1 Raw pixel space
In short-term FFS, methods in raw pixel space have achieved good results.
Recurrent networks.
The exploration of recurrent neural networks in video synthesis is pioneered by PredNet [71], which draws inspiration from predictive coding in neuroscience and employs a recurrent convolutional network for the effective processing of video features. Building on this, PredRNN [110] introduces significant enhancements by modifying Long Short-term Memory (LSTM) with a dual memory structure, aiming for enhanced spatiotemporal modeling. Despite its advancements, it encounters challenges with gradient vanishing in video synthesis tasks. Addressing such limitations, ConvLSTM [85] emerges as a pivotal model, ingeniously integrating LSTM with Convolutional Neural Network (CNN) to proficiently capture motion and spatiotemporal dynamics, a development that has significantly influenced subsequent video synthesis models. Further advancing the field, E3d-LSTM [108] innovatively incorporates 3D convolutions into RNNs and introduces a gate-controlled self-attention module, thereby markedly improving long-term synthesis capabilities. MSPred [97] proposes a hierarchical convolutional and recurrent network operating at multiple temporal frequencies to predict future video frames, as well as other representations such as poses or semantics.
Convolutional networks.
CNNs have been instrumental in the evolution of video synthesis technology. Beginning with GDL [74], the field has seen significant advancements. Following this, PredCNN [120] establishes a new benchmark by outperforming its predecessor PredRNN [110] across various datasets. SDC-Net [80] introduces a novel approach by utilizing a high-resolution video frame synthesis technique that effectively leverages past frames and optical flow. Building on these innovations, the recent introduction of SimVP [35] marks another milestone. This approach revisits the advancements made by ViT [28] and introduces a simplified CNN network, demonstrating that such a configuration can achieve comparable performance in video synthesis. This progression underscores the rapid advancements in convolutional network applications in video synthesis, highlighting a continuous trend toward more efficient and effective models.
Optical-flow-based synthesis.
The optical flow describes the motion of pixels between frames and can be used in flowing pixels of the current frame to synthesize near future frames [70]. Focusing on the enhancement of synthesis quality, FVS [116] utilizes a comprehensive approach by incorporating supplementary information such as semantic maps, instance maps, and optical flow from input frame sequences. This method, while effective, introduces challenges due to increased data modalities and computational demands. OPT [117] estimates the optical flow of video motions in an optimization manner. By continuously refining the current optical flow estimation, the image quality of the next frame can be significantly improved. DMVFN [50] enhances the dense voxel flow [70] estimation by dynamically changing network architecture based on motion magnitude. The study of optical flow estimation is a hot topic [94, 51, 91, 27] in itself. We believe that the near future frames can be synthesized increasingly better under the constantly improving optical flow method. And it is still a challenge to integrate optical flow methods into the goal of improving image synthesis quality [121, 52] and long-term video generation [66].
4.2 Feature space
Synthesizing in raw pixel space often overburdens models by requiring them to reconstruct images from scratch, a task particularly challenging for high-resolution video datasets. This realization has led to a shift in focus among researchers. Rather than grappling with the complexities of pixel-level synthesis, several studies have pivoted towards high-level feature synthesis in feature space, such as segmentation and depth maps. These approaches offer a more efficient way of handling the intricacies of videos [76].
Future semantic segmentation.
Future semantic segmentation represents a progressive approach of video synthesis, primarily focusing on synthesizing semantic maps for forthcoming video frames. This methodology diverges from traditional raw pixel forecasting, turning to semantic maps to narrow the synthesis scope and enrich scene comprehension. In this context, the S2S model [72] stands as a groundbreaking end-to-end system. It processes RGB frames alongside their semantic maps, both as input and output. This integration not only advances future semantic segmentation but also elevates the task of video frame prediction, showcasing the distinct advantage of semantic-level forecasting. Building on this foundation, SADM [7] further innovates by amalgamating optical flow with semantic maps. This fusion leverages optical flow for motion tracking and semantic maps for appearance detailing, employing the former to warp input images and the latter to inpaint occluded areas.
Future depth prediction.
Depth maps, as a 2D data structure containing 3D information, can provide the model with an enhanced perception of the 3D world at minimal cost. Leveraging predictions of future depth maps can assist in future frame synthesis tasks. MAL [68] introduces a meta-learning framework with a two-branch architecture containing future depth prediction and an auxiliary task of image reconstruction. The proposed meta-learning framework improves the quality of synthesized future frames, especially in complex dynamic scenes.
5 Stochastic synthesis
In the early stages, video synthesis is primarily perceived as a low-level computer vision task, with a focus on employing deterministic algorithms to enhance pixel-level metrics such as MSE, PSNR, and SSIM. This approach, however, inherently limits the potential for creative output from these models by confining possible motion outcomes to a single, fixed result [76]. This often results in high pixel-level scores but at the cost of producing blurry images, which significantly diminishes the practical applicability of these algorithms. Recognizing this issue, there has been a paradigm shift in the field of video synthesis, moving away from the reliance on short-term deterministic prediction towards embracing long-term stochastic generation. This transition acknowledges that while stochastic synthesis may yield results that diverge significantly from the ground truth, it plays a crucial role in fostering a more comprehensive understanding and enhancing creativity concerning the evolution of video content.
5.1 Stochasticity modelling
Modeling the uncertain object motion can be achieved by introducing stochastic distributions into deterministic models, or directly leveraging probabilistic models.
5.1.1 Stochastic distributions
In the early stage, VPN [56] uses CNNs for multiple predictions in videos based on pixel distributions and SV2P [4] enhances an action-conditioned model [31] with stochastic distribution estimation for videos. Shifting the focus to a more holistic view of video elements, PFP model [48] presents a probabilistic method for synthesizing semantic segmentation, depth map, and optical flow in videos simultaneously. Additionally, SRVP [33] utilizes Ordinary Differential Equations (ODEs), while PhyDNet [41] employs Partial Differential Equations (PDEs) to calculate stochastic distributions.
5.1.2 Probabilistic models
With the pioneering work of GDL [74] establishing the foundation, adversarial training has significantly advanced future frame synthesis task in predicting uncertain object motions. Similarly, vRNN [16] and GHVAE [114] enhance VAEs through likelihood networks and hierarchical structures respectively, thereby offering another dimension to the ongoing evolution of stochastic synthesis methodologies.
Recognizing that object motion is largely deterministic barring unforeseen events like collisions, SVG [24] models trajectory uncertainty using fixed and learnable priors, which effectively blends deterministic and probabilistic approaches. In a similar vein, but with a focus on enhancing temporal aspects, Retrospective Cycle GAN [62] introduces a sequence discriminator designed to detect fake frames. This idea of scrutinizing frame authenticity is further extended in DIGAN [129], where the emphasis shifts to a motion discriminator that concentrates on identifying unnatural motions.
To overcome pixel-level synthesis challenges in stochastic models, several works introduce intermediate representations. S2S [72] and Vid2Vid [102] integrate adversarial training with future semantic segmentation. Additionally, TPK model [101] leverages a VAE to extract human pose information, and then a GAN for predicting future poses and frames.
It is worth noting that directly modeling stochastic distributions is prone to covering a broader predictive distribution but suffers from poor visual effects. In contrast, probabilistic models are capable of producing sharper results, yet they grapple with issues such as mode collapsing, training difficulties, and substantial computational overhead. Bridging these two methodologies, SAVP [63] combines stochastic distributions with adversarial training aiming to achieve broader predictive distributions with promising visual quality.
5.2 Disentangling components
Stochastic synthesis algorithms primarily focus on the randomness in motion. This approach, however, often neglects the birth-and-death phenomena of objects in videos. Consequently, numerous studies isolate motion from other video elements or artificially manipulate its evolution, providing a clearer understanding of motion dynamics while simplifying the complexity of real-world scenarios.
5.2.1 Content and motion
Video synthesis algorithms confront the challenging complexity of natural video sequences by emphasizing intricate image details. To this end, they focus on effectively modeling appearances through detailed local information, while simultaneously developing a comprehensive understanding of the dynamic global content inherent in videos. However, in applications such as robot navigation and autonomous driving, the importance of understanding object motion patterns supersedes the pursuit of visual aesthetics. This priority shift has encouraged the development of algorithms that emphasize predicting the movement of objects and distinguishing motion from appearance in videos. Early work CDNA [31] has set a precedent by explicitly predicting the motion of objects. It maintains invariant appearance characteristics, which is instrumental in applying the model to objects beyond those encountered during training.
MoCoGAN [95] automatically learns to disentangle motion from content in an unsupervised manner, while the approach of leveraging separate content and motion encoder pathways has also been widely used in various video prediction models. This idea of separating content and motion is further explored in LMC [64], which makes the motion encoder focused on motion prediction based on residual frames, while the content encoder extracts content features from the input frame sequence. MMVP [133] takes a different approach by employing only one image encoder to extract information, and then a two-stream network before the image decoder to separately handle motion prediction and appearance maintenance.
Addressing the stochastic nature of motion, AMC-GAN [53] models multiple plausible outcomes via adversarial training. Transitioning to a different approach, SLAMP [3] adopts a non-adversarial approach but focuses on learning stochastic variables for separate content and motion. Further advancing this field, LEO [111] and D-VDM [84] leverage diffusion models for more realistic disentanglement of content and motion, showcasing the latest advancements in this direction.
5.2.2 Foreground and background
In the process of predicting future frames, the motion dynamics of the foreground (objects) and background (scene) typically exhibit significant differences. Foreground objects typically exhibit more intense movement, whereas scenes tend to remain relatively static. This distinction has steered research towards predicting the motions of these elements separately, offering a more nuanced understanding of video dynamics. A pivotal contribution in this area is DrNet [25], which specifically addresses scenarios where the background remains largely unchanged across video frames. This model ingeniously decomposes images into the content and pose of objects, then harnesses adversarial training techniques to develop a scene discriminator that evaluates whether two pose vectors belong to the same video sequence. Similarly, OCVP-VP [98] utilizes a slot-wise scene parsing network SAVi [58] to achieve segmentation from the scene to the object level. Focusing on such videos, prediction models can streamline the prediction process by eliminating the need to learn complex scene dynamics. Human-centric tasks, such as predicting human movement and interaction with the environment, and object-centric tasks, such as tracking object motion and positioning, both benefit from this approach.
Human-centric.
Future frame synthesis often centers around foreground motion, particularly when it involves intricate human movements. In such scenarios, a common assumption across various specialized datasets is the relatively static nature of the background, which is a typical characteristic in datasets focusing on detailed human motion. This has led to a significant focus in research on understanding and forecasting human poses to enhance the prediction of foreground motion. An example of this approach is seen in the work of DVGPC [13], which innovatively predicts skeleton motion sequences before transforming these sequences into pixel space using a skeleton-to-image transformer. Their method is an effective solution for linking the abstract representation of motion with the practical aspect of video prediction, demonstrating a nuanced understanding of the complexities involved in human-centric FFS tasks.
Object-centric.
The field of object-centric video prediction is first introduced in the work of CVP [124], which lays the foundation for this specialized area of video prediction. SlotFormer [118] introduces transformer-based autoregressive models to learn representations of each object in video sequences. This innovation ensures consistent and accurate tracking of each object over time. The OKID model [21] represents a recent advancement. It uniquely decomposes videos into distinct elements, specifically the attributes and trajectory dynamics of moving objects, employing a Koopman operator. This approach highlights a more detailed method of analyzing object motion within video sequences, differentiating it from previous methodologies.
General.
Methods that concentrate on human poses or objects have demonstrated considerable promise in specific video datasets, yet they encounter limitations due to their reliance on pre-defined structures and struggle with adapting to variable backgrounds, which hampers their ability to generalize. This challenge is evident in their performance, which, while effective under certain conditions, falters when faced with dynamic background shifts, revealing a lack of the versatility required for wider applications. Bridging this gap, MOSO [92] emerges as a notable approach, identifying motion, scene, and object as the pivotal elements of a video. It delves deeper into content analysis by distinguishing between scene and object. Scene and object are considered as a further breakdown of content, where scene represents the background and object represents the foreground. MOSO’s innovative contribution lies in its two-stage network tailored for general video analysis. Initially, the MOSO-VQVAE model dissects video frames into token-level representations, honing its capabilities through a video reconstruction task. Subsequently, the model employs transformers in its second stage, tackling disparities in masked tokens. This strategic design equips the model to handle a variety of tasks at the token level, including video prediction, interpolation, and unconditional video generation.
5.3 Motion-controllable synthesis
In the field of future frame synthesis, one specialized research direction has emerged that focuses on the explicit control of motion. This approach is distinct in its emphasis on forecasting future object positions based on user-defined instructions, diverging from the conventional reliance on past motion trends. The central challenge in this domain lies in the synthesis of videos that adhere to these direct instructions while maintaining a natural and coherent flow, a task that demands a nuanced understanding of both user intent and the dynamics of motion within a video context. This challenge highlights the intricate balance between user control and automated imagination, marking a significant shift in how future frame synthesis models are conceptualized and implemented.
Strokes.
There is no historical motion information to be used for video generation from one still image, hence several methods have emerged allowing for interactive user control. iPOKE [8] introduces techniques where local interactive strokes and pokes enable users to deform objects in one still image to generate a sequence of video frames. These strokes indicate the user’s intended motion for the objects. Following this innovative pathway, Controllable-Cinemagraphs model [73] introduces a method to interactively control the animation of fluid elements. The advancements in the field underscore the growing importance of user-centric approaches in the realm of motion-controllable future frame synthesis.
Instructions.
The integration of instructions of various modalities, including local strokes, sketches, and texts, is increasingly common in works aiming to capture user-specified motion trends. VideoComposer [104] synthesizes videos by combining text descriptions, hand-drawn strokes, and sketches. This approach respects textual, spatial, and temporal constraints, leveraging video latent diffusion models and motion vectors for explicit dynamic guidance. Essentially, It can generate videos that align with user-defined motion strokes and shape sketches. In a similar vein, DragNUWA [125] primarily leverages text for video content description and strokes for future motion control, enabling the generation of customizable videos. These attempts advance the field of video generation by broadening the range of user input modalities.
6 Generative synthesis
In video analysis, the focus shifts towards algorithms designed for generative video synthesis, especially when dealing with videos that exhibit stochastic birth-and-death events. These events introduce unpredictability as objects come into existence and vanish. These algorithms necessitate a profound understanding of the underlying physical principles governing the real world to tackle such complexities. Instead of relying on simplistic linear motion predictions extrapolated from historical frames, they embrace the challenge with sophisticated and imaginative modeling techniques. As a result, tasks like transforming a single static image into a dynamic video, often referred to as the image animation problem, emerge as promising candidates for applying generative video prediction techniques.
6.1 Diffusion-based generation
Diffusion models [47] have become dominant in the image generation realm. The Latent Diffusion Model (LDM) [81] explores the generation capability in the latent space of images, significantly improving computational efficiency and reducing costs. LDM has paved the way for introducing diffusion models to the video generation realm. The extensions of LDM exhibit robust generative capabilities [100]. Furthermore, Video LDM [9] leverages pre-trained image models for video generation, offering multi-modal, high-resolution, long-term video generation ability. Similarly, SEINE [20] introduces a versatile video diffusion model, enabling the creation of transition sequences to generate longer videos from shorter clips.
Recently, there has been a surge in research efforts aimed at harnessing additional information in conjunction with RGB images to fulfill the text-guided video completion task. LFDM [75] builds upon the latent diffusion models to synthesize optical flow sequences in latent space based on textual guidance. Seer [40] inflates Stable Diffusion [81] along the temporal axis, enabling the model to utilize natural language instructions and reference frames to imagine multiple variations of future outcomes. Emu Video [37] takes a different approach by initially generating an image conditioned on textual guidance and subsequently extrapolating it into a video. Consequently, this method can be flexibly applied to create video based on a given image with different textual inputs. DynamiCrafter [119] broadens the application of text-guided image animation to open-domain images. SparseCtrl [42] allows for sketch-to-video generation, depth-to-video generation, and video prediction with an expanded input range.
Some works attempt to provide a more precise understanding of the guidance of text and preserve the guidance information along the time dimension. MicroCinema [107] adopts a divide-and-conquer strategy to tackle appearance and temporal coherence challenges. This model employs a two-stage generation pipeline, which initially utilizes any existing text-to-image generator for the creation of the initial image, and subsequently introduces a dedicated text-guided video generation framework specifically designed for motion modeling. LivePhoto [19] introduces a framework that incorporates motion intensity as a supplementary factor to enhance control over desired motions. Additionally, it proposes a text re-weighting mechanism to emphasize motion descriptions, showcasing impressive performance on text-guided video synthesis tasks. I2VGen-XL [131] utilizes static images for semantic and qualitative guidance, showcasing the diverse approaches in text-guided video synthesis research.
Although diffusion models have almost become the most popular generative models, many researchers are attempting to preserve the architecture of autoregressive models to simultaneously utilize the advantages of both mechanisms [113]. Despite the limitations in data availability and model scalability that hindered the early generative video synthesis algorithms mentioned above, they laid the groundwork for integrating LLMs and diffusion models in the field of video generation. Recently, GAIA-1 [49] and Sora [11] leverage the strengths of both diffusion models and large language models to achieve more creative, general, and scalable video synthesis.
6.2 Token-based generation
In the field of image/video generation, while diffusion-based methods attract the largest community and most researchers, these algorithms generally have much smaller model sizes compared to the large-scale language models (LLMs) commonly seen today. As a result, many studies focus on exploring how transformers can accomplish these tasks. The advantage of this approach lies in leveraging a large amount of related research and optimization methods from LLMs to validate the scaling laws in the visual domain. Implementing transformers for future frame synthesis involves two key components: an efficiently scalable LLM framework, and a good image/video tokenizer. Innovations like VQ-VAE [96] and VQGAN [30] combine autoregressive models and adversarial training strategies to tackle image quantization and tokenization. However, an effective visual tokenizer should minimize the number of tokens per image or video clip segment while ensuring near-lossless visual reconstruction. During LLM training, the typical context size ranges from 8k to 32k tokens, while lossless reconstruction of a $256\times 256$ resolution image usually requires no fewer than 500 tokens with extremely large vocabulary size [127], with high-resolution images demanding thousands of tokens. This makes it difficult for the model to process long sequences of video clips during training, severely limiting the model’s video generation capabilities.
Before the rise of LLMs [2], the transformative impact of transformers has been evident in time-series modeling. Video Transformer [112] applies transformers to create an autoregressive model, pioneering the application of transformer architectures in the context of video synthesis. While this autoregressive framework marked a successful attempt in the realm of natural video generation, it also inherited the consistent drawbacks of transformers, namely extensive training resource consumption and prolonged inference time. The Latent Video Transformer (LVT) [78] proposes a novel approach by considering the modeling of dynamics and predicting future features in an autoregressive manner within a latent space to release the computation burden. The NUWA [115] framework proposes a versatile 3D transformer encoder-decoder architecture, adaptable to various data modalities and tasks, further highlighting the potential of transformers in video synthesis. NUWA-Infinity [67] expands on this with an innovative generation mechanism, with the ambitious goal of achieving infinite high-resolution video generation. This progress signifies a continuous effort to unify generative tasks across different modalities, showcasing the evolving landscape of future frame synthesis research.
LVM [5] introduces sequential modeling to expand the learning capabilities of large-scale vision models, showcasing the scalability and flexibility of sequence models in handling in-context learning. The paper also proposes a concept of ”visual sentence” where a sequence of images with intrinsic relationships is arranged similarly to a sentence in language, allowing the model to leverage the information within the sequence to perform sentence continuation. This unified manner enables the model to utilize the prompt information of the visual sentence to synthesize future frames or complete other visual tasks without relying on any non-pixel-level knowledge. The in-context learning is not unfamiliar in the visual domain. Painter [103] introduces a general framework for visual learning that enables images to ”speak” through in-context visual learning, enhancing both image generation and understanding capabilities. Similarly, SegGPT [105] investigates using a GPT architecture for image segmentation and proposes the concept of segmenting everything demonstrating the possibility of achieving unified image segmentation in an unsupervised learning environment, further advancing generalization in visual segmentation tasks.
In the video generation domain, MAGVIT [127] presents a masked generative video Transformer that efficiently processes video by masking certain parts and generating the missing segments. Building on this, MAGVIT-v2 [128] proposes that language models can outperform diffusion models in visual generation tasks, emphasizing the critical role of visual tokenizers in visual generation, which offers new insights into the future development of visual generation tasks. Utilizing the powerful visual tokenizer proposed by MAGVIT-v2, VideoPoet [59] introduces a large language model for zero-shot video generation, pushing the boundaries of unsupervised learning in the video generation field. It allows users to generate or modify videos based on high-level textual prompts. The model excels in capturing the temporal and contextual relationships within video data, enabling it to perform tasks such as video generation, continuation, and editing directly from text descriptions. These studies collectively integrate visual tokenizers and large language models to offer unified and scalable frameworks for visual learning, driving breakthroughs in future frame synthesis tasks.
As mentioned above, text-guided generative video synthesis algorithms are designed to create a sequence of frames by integrating context frames and textual guidance. Text-guided Video Completion (TVC) task involves completing videos based on various conditions, including the first frame (video prediction task), the last frame (video rewind task), or a combination of both (video transition task), all of which are guided by textual instructions. MMVG [34] addresses the TVC task by utilizing autoregressive encoder-decoder architectures, integrating text and frame features, resulting in a unified framework capable of handling multiple video synthesis tasks.
7 Future research
Based on the overall development trends in the fields of Artificial Intelligence and Computer Vision, and considering the challenges in the future frame synthesis domain discussed in Section 2.2, we contend that the trajectory of video synthesis research should diverge from the traditional trends of data simplification and model lightweight in low-level tasks. The imperative should be to harness substantial computational resources and high-resolution, long-duration video data. Future research should prioritize the formulation of evaluation metrics that incentivize stochastic synthesis, thus expanding the potential of simulation of the human world. The ultimate objective is to cultivate video synthesis models with a profound comprehension of the inherent dynamics within videos. Such models could execute video generation over extended time horizons with high stochastic complexity in the real world.
Dataset Year Category # Videos # Clip Frames Resolution Extra Annotations
KTH Action [83] 2004 Human 2, 391 95 160 $\times$ 120 Class
Caltech Pedestrian [26] 2009 Human 137 1, 824 640 $\times$ 480 Bounding Box
HMDB51 [60] 2011 Human 6, 766 93 414 $\times$ 404$\dagger$ Class
UCF101 [89] 2012 Human 13, 320 187 320 $\times$ 240 Class
J-HMDB [54] 2013 Human 928 34 320 $\times$ 240 OF, Ins, HJ, Class
KITTI [36] 2013 Traffic 151 323 1242 $\times$ 375 OF, BBox, Sem, Ins, Depth
Penn Action [132] 2013 Human 2, 326 70 480 $\times$ 270$\dagger$ Human Joint, Class
SJTU 4K [88] 2013 General 15 300 3840 $\times$ 2160 -
Sports-1M [57] 2014 Human 1, 133, 158 variable variable Class
Moving MNIST [90] 2015 Simulation 10, 000 20 64 $\times$ 64 -
Cityscapes [22] 2016 Traffic 46 869 2048 $\times$ 1024 Semantic, Instance, Depth
YouTube-8M [1] 2016 General 8, 200, 000 variable variable Class
Robotic Pushing [31] 2016 Robot 59, 000 25 640 $\times$ 512 Class
DAVIS17 [77] 2017 General 150 73 3840 $\times$ 2026$\dagger$ Semantic
Something-Something [38] 2017 Object 220, 847 45 427 $\times$ 240$\dagger$ Text
ShapeStacks [39] 2018 Simulation 36, 000 16 224 $\times$ 224 Semantic
SM-MNIST [24] 2018 Simulation customize customize 64 $\times$ 64 -
D2-City [17] 2019 Traffic 11, 211 750 1080p / 720p BBox
Kinetics-700 [15, 14] 2019 Human 650, 000 250 variable Class
RoboNet [23] 2019 Robot 161, 000 93 64 $\times$ 48 -
Vimeo-90K [121] 2019 General 91, 701 7 448 $\times$ 256 -
BDD100K [126] 2020 Traffic 100, 000 1175 1280 $\times$ 720 BBox, Semantic, Depth
nuScenes [12] 2020 Traffic 1, 000 40 1600 $\times$ 900 BBox, Semantic
WebVid [6] 2021 General 10, 732, 607 449 596 $\times$ 336 Text
X4K1000FPS [87] 2021 General 4, 408 65 4096 $\times$ 2160 -
SportsSlomo [18] 2024 Human 130,000 7 $1280\times 720$
InternVideo2 [109] 2024 General 2, 000, 000 variable variable Text, Action
OpenDV-YouTube [123] 2024 Traffic 2, 139 100,000+ variable Text
denotes the mean value. $\dagger$ denotes the median value.
Table 1: Summary of the most used video prediction datasets, including the total number of videos, frame count for each video clip, image resolution, and additional annotations, etc. (OF: Optical Flow, BBox: Bounding Box, Sem: Semantic, Ins: Instance, HJ: Human Joints)
8 Conclusion
In this survey, we have discussed various aspects of future frame synthesis, covering key topics such as predominant datasets, evolving algorithms, current challenges, and future trends. We propose a novel taxonomy based on algorithm stochasticity, showcasing a significant shift from deterministic approaches toward generative methodologies. This survey emphasizes the importance of balancing pixel-level accuracy with a deep understanding of complex scene dynamics in video synthesis. Additionally, we delve into the intricacies of stochastic birth-and-death phenomena, advocating for enhanced evaluation metrics and the utilization of substantial computational resources and large-scale video datasets. These insights aim to guide future research directions in video synthesis. As the field advances, we anticipate the development of models that provide a more profound and nuanced grasp of nature in the real world. This evolution promises to enhance accuracy, efficiency, and creativity, paving the way for novel applications and research.
Acknowledgments
This work is supported by Beijing Natural Science Foundation L231014.
Appendix
9 Datasets
We offer an overview of the most widely used datasets in video synthesis, highlighting their amounts of data and additional supervisory modalities, thereby providing a comprehensive picture of the current dataset landscape in this field.
10 Application realms
The applications of future frame synthesis span a wide range of domains, showcasing its significance in different fields.
World Model.
World models [43, 134] provide a more general framework for simulating and predicting the behavior of complex systems. A world model is often used in reinforcement learning and robotics to enable agents to make informed decisions and take actions that lead to desired outcomes. To construct a world model, future frame synthesis is an important learning objective [44, 45, 106].
Autonomous driving.
Future frame synthesis is vital for self-driving cars and drones to anticipate the movement of objects, pedestrians, and other vehicles, ensuring safe navigation. GAIA-1 [49] employs a unified world model, combining multi-modal large language models and diffusion procedures to predict control signals and future frames.
Robot navigation.
Future frame synthesis is employed in guiding robots through dynamic environments, allowing them to plan their paths, move objects, and avoid obstacles effectively [32].
Cinema industry.
Future frame synthesis is utilized in special effects, animation, and pre-visualization, aiding filmmakers in creating realistic scenes and enhancing the overall cinematic experience [73].
Meteorological community.
Future frame synthesis plays a crucial role in weather forecasting, helping meteorologists simulate and predict atmospheric conditions, contributing to improved accuracy in weather predictions [86].